# Minghan Chen  Guikun Chen  Wenguan Wang  Yi Yang

### v1

@@ Abstract
DETR introduces a simplified one-stage framework for scene graph generation (SGG). However, DETR-based SGG models face two challenges: i) Sparse supervision, as each image typically contains fewer than 10 relation annotations, while the models employ over 100 relation queries. This sparsity arises because each ground truth relation is assigned to only one single query during training. ii) False negative samples, since one ground truth relation may have multiple queries with similar matching scores. These suboptimally matched queries are simply treated as negative samples, causing the loss of valuable supervisory signals. As a response, we devise Hydra-SGG [^1] Hydra is an abbreviation for “Hybrid Relation Assignment”. The name is chosen because it reflects the multi-branch structure of our model, which resembles the multiple heads of the Hydra in Greek mythology., a one-stage SGG method that adopts a new Hybrid Relation Assignment. This assignment combines a One-to-One Relation Assignment with a newly introduced IoU-based One-to-Many Relation Assignment. Specifically, each ground truth is assigned to multiple relation queries with high IoU subject-object boxes. This Hybrid Relation Assignment increases the number of positive training samples, alleviating sparse supervision. Moreover, we, for the first time, empirically show that self-attention over relation queries helps reduce duplicated relation predictions. We, therefore, propose Hydra Branch, a parameter-sharing auxiliary decoder without a self-attention layer. This design promotes One-to-Many Relation Assignment by enabling different queries to predict the same relation. Hydra-SGG achieves state-of-the-art performance with 10.6 mR@20 and 16.0 mR@50 on VG150, while only requiring 12 training epochs. It also sets a new state-of-the-art on Open Images V6 and GQA.
@@ 1 Introduction
Figure 1: Comparisons with other SGG methods in terms of mR@50 and training epochs on VG150 [68].
@@
Scene graph is a data structure that describes the entities (objects) in a scene and the relations between these objects [3]. Scene graph generation (SGG) has attracted significant research attention [20; 47; 3; 40] due to its ability to enhance machines’ semantic comprehension of visual content. It has been widely adopted in various downstream applications, such as robotic vision and interaction [44; 55; 73; 24], image synthesis and manipulation [72; 9; 22], visual question answering [21; 61; 38], and video understanding [78; 71; 63; 69].
@@
Mainstream SGG methods [77; 59; 47; 83; 56; 30] work in a two-stage fashion. First, an off-the-shelf object detector extracts all entities within an image. Then, the extracted entities are permuted, yielding $N(N\!-\!1)$ entity pairs for $N$ detected entities. These entity pairs are used to predict the relationships between the corresponding entities. However, the two-stage methods face a critical limitation: they predict relations for all entity pairs, even though many pairs do not participate in any relations. This incurs heavy computational overhead and time-consuming inference in complex scenes with numerous entities.
@@
Figure 2: GT stands for Ground Truth. (a) Previous DETR-based SGG methods such as RelTR [6] and SGTR [37] match each GT relation with only one query. (b) Our Hybrid Relation Assignment utilizes both One-to-One and One-to-Many assignments, generating more positive samples, and thus accelerating training.
@@
Recently, the inherent simplicity of DETR-based SGG methods has led to an ongoing paradigm shift apart from the two-stage framework. Specifically, a sequence of visual tokens, mapped from the input image, interacts with a predefined set of relation queries in a Transformer decoder for simultaneous object detection and relation prediction. Then, One-to-One set matching (i.e., Hungarian matching [29]) is used to assign ground truth labels to the predictions (${}_{\!}$ 2a). This set-prediction framework allows DETR-based SGG models to eliminate hand-designed components, such as Non-Maximum Suppression (NMS), which are commonly used in traditional two-stage methods.
@@
Unfortunately, one significant drawback of DETR-based SGG models is their slow convergence. For instance, one-stage RelTR [6] requires 150 training epochs to converge, while a recent two-stage method, PE-Net [83], only needs 32 epochs. This drawback can be attributed to the sparse relation supervision induced by Hungarian matching, which assigns each ground truth to only one relation query. The sparsity of relation annotations per image further exacerbates this issue. Taking VG150 [68] train as an example, there are, on average, only 5.5 ground truth relation triplets per image [40]. This means that each image provides a mere 5.5 positive relation queries to optimize the loss functions. In the case of RelTR [6], which has 200 predefined relation queries, only $2.75\%$ of these queries are positive samples per optimization step. Consequently, relation queries require more optimization steps to learn due to the limited number of positive samples for training. Furthermore, approximately $50\%$ [27] of the plausible but suboptimally matched queries (i.e., queries with correct subject-object classification and IoU $>$ 0.6 for both boxes) are simply treated as no-relation due to the One-to-One constraint. This constraint discards valuable supervisory signals by treating suboptimal yet informative queries as negative samples. These queries, while not the best matches for ground truth labels, may capture informative relational cues that could contribute to model learning. Simply assigning these queries as negative samples may introduce false negatives, which in turn leads to label noise and performance degradation [42].
To improve the training efficacy and performance of DETR-based SGG models, we introduce Hydra-SGG. The core of Hydra-SGG is Hybrid Relation Assignment (${}_{\!}$ 2b); it synergizes One-to-One and One-to-Many Relation Assignment strategies, providing over 50% more positive queries per training step than previous arts such as RelTR [6]. Specifically, an IoU-based One-to-Many Relation Assignment is devised, allowing each ground truth label to be assigned to multiple relation queries with high IoU between their predicted subject-object boxes and the ground truth. Moreover, we provide empirical evidence that removing self-attention in the decoder leads to different queries making duplicate relation predictions. With such insight, we introduce an auxiliary branch, dubbed Hydra Branch, to further enhance One-to-Many Assignment. This branch is a decoder without self-attention layers and shares all other parameters with the original decoder. Hydra Branch encourages the predictions of different queries to converge to the same relation, creating a synergy with the proposed One-to-Many Relation Assignment. This synergistic integration of Hydra Branch and One-to-Many Relation Assignment further improves the overall performance (§3.3).
Hydra-SGG makes three key contributions to the field of SGG: First, the proposed Hybrid Relation Assignment increases both the quality and quantity of positive relation queries. By allowing each ground truth to be assigned to multiple queries based on their IoU scores, Hydra-SGG effectively alleviates the sparse supervision issue. Second, as a one-stage SGG model, Hydra-SGG not only achieves state-of-the-art results on the challenging VG150 [68] but also exhibits a remarkably fast convergence rate. It converges faster than existing one-stage SGG counterparts [6; 37; 18] with 10$\times$ fewer training epochs. Third, Hydra Branch does not introduce any architectural modifications or additional parameters. It is only used during training and is dropped in the inference stage, causing no additional computational load or delay in the inference stage.
@@
Extensive experiments on three challenging SGG benchmarks, VG150 [68], GQA [16], and Open Images V6 [31], demonstrate the effectiveness of our proposed method. Notably, Hydra-SGG achieves 16.0 mR@50 on VG150 test in only 12 epochs (${}_{\!}$ 1), surpassing the previous state-of-the-art one-stage method SGTR [37] by +4.0 and two-stage method PE-Net [83] by +3.6.
@@ 2 Related Work
Two-stage SGG. Lu et al. [47] first proposed SGG task and designed a two-stage pipeline. Later, many SGG models have been built on it, using various techniques such as recurrent neural networks or its variants [77; 59; 68; 65], visual translation embedding [80; 17], graph neural networks [75; 57; 67; 75; 70], external or internal knowledge integration [5; 11; 76; 1; 39; 74], attention mechanisms [23; 50; 8], and Transformer-based models [48; 6; 37; 56; 7; 30; 54].
@@
Despite promising performance, two-stage SGG methods rely heavily on manually designed modules, such as NMS, anchor generation, and entity pairing generation modules. In addition, their designs typically involve separate stages for detection, pairing, and relation classification, resulting in a complicated pipeline that cannot be trained in a fully end-to-end manner. Furthermore, these methods predict dense entity pairs in inference, leading to high time complexity and computational burden. For example, if there are 100 detected entities, approximately 10,000 pairs need to be processed. In contrast, this paper proposes Hydra-SGG, a one-stage SGG method that offers significant advantages over two-stage methods. Hydra-SGG enables end-to-end training and surpasses the performance of two-stage methods.
@@
One-stage SGG. One-stage methods have gained increasing attention for their simplicity and end-to-end training ability. Early works in this line adopt CNN-based one-stage detectors [43] or query-based sparse R-CNN [62] for direct relationship prediction. Recently, the DETR [2] framework significantly advances SGG and Human-Object Interaction [6; 37; 58; 86; 41; 26; 79; 27; 18]. This framework enables end-to-end training by associating ground truth labels with output queries. For instance, SGTR [37] uses predicate and entity queries to associate relation features, while RelTR [6] employs disentangled queries to extract relation features and directly predict relations. SpeaQ [27] designs multiple groups of relation-specific queries to learn relation representations effectively.
DETR-based SGG models eliminate NMS but suffer from slow convergence due to sparse relation supervision, requiring significantly more training epochs than two-stage counterparts (e.g., 150 for RelTR [6] vs. 32 for PE-Net [83]). To address this limitation, we propose Hydra-SGG, which introduces a Hybrid Relation Assignment. This method exhibits fast convergence compared to two-stage approaches like PE-Net [83] and VETO [56]. Like other DETR-based SGG models, Hydra-SGG eliminates the need for complicated pipelines or hand-crafted components such as NMS.
@@
DETR for Object Detection. The introduction of DETR by Carion et al. [2] marks a significant shift away from traditional CNN-based object detection models [53; 51; 52; 42; 14], adopting an end-to-end trainable approach with a novel application of the Transformer architecture and bipartite matching. While DETR streamlines the detection process, it requires 500 epochs to converge [84]. This spurs a wave of research that focuses on improving low training efficacy, including enhancing training signal [4; 15; 19; 85], the adoption of anchor boxes [45; 49; 66], and the leverage of efficient attention mechanism [84; 10]. For instance, Co-DETR [85] introduces extra groups of object queries and DN-DETR [32] utilizes noisy queries to increase positive samples.
@@ 3 Hydra-SGG
§3.1 introduces our baseline model, which employs One-to-One Relation Assignment. Next, §3.2 describes Vanilla Hydra-SGG, which enhances the baseline by introducing a Hybrid Relation Assignment strategy. This strategy combines the standard One-to-One Relation Assignment with our One-to-Many Relation Assignment, bringing more than $50\%$ positive samples to each training step compared to the baseline. Finally, §3.3 details the complete Hydra-SGG model, which includes Hydra Branch, an auxiliary decoder that promotes One-to-Many Relation Assignment.
@@ 3.1 One-to-One Relation Assignment Baseline
Baseline Architecture. Our baseline model is composed of a Backbone, a Encoder, and a RelDecoder (Relation Decoder). The relation queries interact with image tokens extracted from the input image by Backbone and Encoder. The updated queries are processed by box regression heads, entity classification heads, and a relation classification head to obtain the final predictions.
@@
  • One-to-One Relation Assignment Loss. In One-to-One Relation Assignment Loss $\mathcal{L}_{\text{o2o}}$ [6; 37], Hungarian matching (HM) is employed to find the optimal correspondence between the predicted relation triplets $\bar{\bm{R}}$ and the ground truth triplets $\bm{R}$. The loss function can be formulated as:
@@
    The classification losses (including entity and relation) and box regression losses are computed between matched predictions and ground truth labels, which is identical with the previous models [6; 37]. The Hungarian matching hyperparameters are the same as RelTR [6]. A detailed implementation of the loss function is provided in Appendix. The evaluation of our baseline model is given in §4.3.
@@ 3.2 Vanilla Hybrid Relation Assignment
As mentioned in §1, VG150 [68] contains an average of only 5.5 ground truth relation triplets per image in train. Our baseline adopts One-to-One Relation Assignment and only assigns approximately 5 out of 300 relation queries to match the ground truth relation triplets for each image (i.e., $N\!=\!300$). This severe sparse supervision reduces the training efficacy, as only about 2% of queries in each training step match with ground truth labels for learning, while the remaining 98% are treated as negative samples. Consequently, the model requires a significantly higher number of training steps to learn effectively, leading to slow convergence.
@@
Figure 4: (a) presents the average number of positive samples of VG150 train and val for One-to-One and Hybrid Relation Assignment. (b) Percentage increase in positive samples achieved by Hybrid Relation Assignment compared to the One-to-One baseline. (c) ADS on VG150 val. (d)-(e) The visualizations show that for the same group of queries that previously predicted different relations, removing the self-attention layers causes them to make the same predictions. The Q ID column represents the ID for each relation query.
@@ 3.3 Complete Hydra-SGG
Each RelDecoder layer contains self- and cross-attention layers, and a point-wise feed-forward network (FFN). The self-attention layer enables inter-query interactions [64], while cross-attention and FFN do not explicitly support query interactions. Previous studies indicate that self-attention helps inhibit duplicate predictions [2; 6]. In this work, we further find that the self-attention layer in RelDecoder is crucial to reduce duplicated relation predictions.

### v2

@@ Abstract
DETR introduces a simplified one-stage framework for scene graph generation (SGG) but faces challenges of sparse supervision and false negative samples. The former occurs because each image typically contains fewer than 10 relation annotations, while DETR-based SGG models employ over 100 relation queries. Each ground truth relation is assigned to only one query during training. The latter arises when one ground truth relation may have multiple queries with similar matching scores, leading to suboptimally matched queries being treated as negative samples. To address these, we propose Hydra-SGG, a one-stage SGG method featuring a Hybrid Relation Assignment. This approach combines a One-to-One Relation Assignment with an IoU-based One-to-Many Relation Assignment, increasing positive training samples and mitigating sparse supervision. In addition, we empirically demonstrate that removing self-attention between relation queries leads to duplicate predictions, which actually benefits the proposed One-to-Many Relation Assignment. With this insight, we introduce Hydra Branch, an auxiliary decoder without self-attention layers, to further enhance One-to-Many Relation Assignment by promoting different queries to make the same relation prediction. Hydra-SGG achieves state-of-the-art performance on multiple datasets, including VG150 (16.0 mR@50), Open Images V6 (50.1 weighted score), and GQA (12.7 mR@50). Our code and pre-trained models will be released on Hydra-SGG.
@@ 1 Introduction
Figure 1: Comparison with other SGG methods in mR@50 and training epochs on VG150 [75].
@@
A scene graph is a data structure that describes the entities (objects) in a scene and the relations between these objects [3]. Scene graph generation (SGG) has attracted significant research attention [20; 50; 3; 43; 71] due to its ability to enhance machines’ semantic comprehension of visual content. It has been widely adopted in various downstream applications, such as robotic vision and interaction [47; 60; 80; 48], image synthesis and manipulation [79; 10; 22], visual question answering [21; 66; 41], and video understanding [78; 68; 76; 81].
@@
Mainstream SGG methods [85; 64; 50; 91; 61; 29; 35] work in a two-stage fashion. First, an off-the-shelf object detector extracts all entities within an image. Then, the extracted entities are permuted, yielding $N(N\!-\!1)$ entity pairs for $N$ detected entities. These entity pairs are used to predict the relationships between the corresponding entities. However, the two-stage methods face a critical limitation: they predict relations for all entity pairs, even though many pairs do not participate in any relations. This incurs heavy computational overhead and time-consuming inference, especially in complex scenes with numerous entities and intricate interactions.
@@
Figure 2: (a) Previous DETR-based SGG methods such as RelTR [7] and SGTR [40] match each GT relation with only one query. (b) Our Hybrid Relation Assignment utilizes both One-to-One and One-to-Many assignments, generating more positive samples and thus accelerating training.
@@
Recently, the simplicity of DETR-based SGG methods [7; 40] has led to an ongoing paradigm shift apart from the two-stage framework. Specifically, a sequence of visual tokens, mapped from the input image, interacts with a predefined set of relation queries in a Transformer decoder for simultaneous object detection and relation prediction. Then, One-to-One set matching (i.e., Hungarian matching [28]) is used to assign ground truth labels to the predictions (${}_{\!}$ 2a). This set-prediction framework allows DETR-based SGG models to eliminate hand-designed components, such as Non-Maximum Suppression (NMS), which are commonly used in traditional two-stage methods.
@@
Unfortunately, one significant drawback of DETR-based SGG models is their slow convergence. For instance, one-stage RelTR [7] requires 150 training epochs to converge, while a recent two-stage method, PE-Net [91], only needs 32 epochs. This drawback can be attributed to the sparse relation supervision induced by Hungarian matching, which assigns each ground truth to only one relation query. The sparsity of relation annotations per image further exacerbates this issue. For instance, VG150 [75] train averages only 5.5 ground truth relation triplets per image [43]. This means that each image provides a mere 5.5 positive relation queries to optimize the loss functions. In the case of RelTR [7], which has 200 predefined relation queries, only $2.75\%$ of these queries are positive samples per optimization step. Consequently, relation queries require more optimization steps to learn due to the limited number of positive samples for training. Furthermore, approximately $50\%$ [26] of the plausible but suboptimally matched queries (i.e., queries with correct subject-object classification and IoU $>$ 0.6 for both boxes) are simply treated as no-relation due to the One-to-One constraint of Hungarian matching. This constraint discards valuable supervisory signals by treating suboptimal yet informative queries as negative samples. These queries, while not the best matches for ground truth labels, may capture informative relational cues that could contribute to model learning. Simply assigning these queries as negative samples may introduce false negatives, which in turn leads to label noise and performance degradation [45].
To accelerate the training of DETR-based SGG models, we introduce Hydra-SGG [^1] Hydra is an abbreviation for “Hybrid Relation Assignment”. The name is chosen because it reflects the multi-branch structure of our model, which resembles the multiple heads of the Hydra in Greek mythology., an efficient framework that addresses the slow convergence problem in one-stage DETR-based SGG models. The cores of Hydra-SGG are Hybrid Relation Assignment and Hydra Branch. Specifically, Hybrid Relation Assignment synergizes One-to-One and One-to-Many Relation Assignment strategies, providing over 50% more positive queries per training step than previous arts such as RelTR [7]. To further enhance this assignment strategy, we introduce an auxiliary branch called Hydra Branch. This branch is specifically designed to encourage different queries to predict duplicated relations by removing self-attention in the decoder, creating a synergistic effect with our Hybrid Relation Assignment. The branch shares all other parameters with the original decoder, and this intentional design choice leads to improved supervision signals during training (§3.3).
Hydra-SGG makes three main contributions to the field of SGG: First, we propose an efficient framework that effectively addresses the slow convergence problem in one-stage DETR-based SGG models through a hybrid query-label assignment and synergistic architectural design. Second, our Hybrid Relation Assignment strategy significantly increases relation supervision signals by mining false negative samples to increase positive samples during training, providing over 50% more positive samples per training step. Third, the proposed Hydra Branch complements our assignment strategy by encouraging duplicate relation predictions during training, while maintaining inference efficiency as it shares parameters with the original decoder and is only used during training. Hydra-SGG achieves state-of-the-art performance with remarkable training efficiency, converging 10$\times$ faster than existing one-stage SGG counterparts [7; 40; 18].
@@
Extensive experiments on three challenging SGG benchmarks, VG150 [75], Open Images V6 [30], and GQA [16], demonstrate the effectiveness of our Hydra-SGG. It achieves 16.0 mR@50 on VG150 test in only 12 epochs (${}_{\!}$ 1), surpassing the previous state-of-the-art one-stage method SGTR [40] by +4.0 and two-stage method PE-Net [91] by +3.6.
@@ 2 Related Work
Two-stage SGG. Lu et al. [50] first proposed SGG task and designed a two-stage pipeline. Later, many SGG models have been built on it, using various techniques such as recurrent neural networks or its variants [85; 64; 75; 70; 58], visual translation embedding [88; 17], graph neural networks [83; 62; 74; 83; 77], external or internal knowledge integration [6; 12; 84; 1; 42; 82; 35; 4; 72], attention mechanisms [23; 53; 9], and Transformer-based models [51; 7; 40; 61; 8; 29; 59].
@@
Despite the promising performance, two-stage SGG methods rely heavily on manually designed modules, such as NMS, anchor generation, and entity pairing generation modules [85; 64; 9]. In addition, their designs typically involve separate stages for detection, pairing, and relation classification, resulting in a complicated pipeline that cannot be trained in a fully end-to-end manner. Furthermore, these methods predict dense entity pairs in inference, leading to high time complexity and computational burden. In contrast, this paper proposes Hydra-SGG, a one-stage SGG method that offers significant advantages over two-stage methods. Hydra-SGG enables end-to-end training and surpasses the performance of two-stage methods.
@@
One-stage SGG. One-stage methods have gained increasing attention for their simplicity and end-to-end training ability. Early works in this line adopt CNN-based one-stage detectors [46] or query-based sparse R-CNN [67] for direct relationship prediction. Recently, the DETR [2] framework significantly advances SGG and Human-Object Interaction [54; 92; 7; 40; 63; 95; 44; 25; 86; 26; 18; 38; 37; 73]. This framework enables end-to-end training by associating ground truth labels with output queries. Existing DETR-based SGG methods focus on different aspects: SGTR [40] and DSGG [13] explore query designs for relation feature extraction, SpeaQ [26] investigates relation-specific query grouping strategies to improve the specialization and discrimination of queries, while other works focus on architectural innovations [7] and lightweight frameworks [18].
However, DETR-based SGG models, while eliminating NMS, face a critical challenge of slow convergence due to sparse relation supervision, requiring significantly more training epochs than two-stage counterparts (e.g., 150 for RelTR [7] vs. 32 for PE-Net [91]). Our proposed Hydra-SGG specifically addresses this limitation through a query-label assignment and architectural design while maintaining the advantages of DETR-based approaches. Hydra-SGG achieves remarkably fast convergence with only 12 epochs, surpassing both one-stage counterparts and two-stage methods to achieve state-of-the-art performance across multiple SGG datasets (16.0 mR@50 on VG150 [75], 50.1 weighted score on Open Images V6 [30], and 12.7 mR@50 on GQA [16]).
@@
DETR for Object Detection. The introduction of DETR by Carion et al. [2] marks a significant shift away from traditional CNN-based object detection models [57; 55; 56; 45; 14], adopting an end-to-end trainable approach with a novel application of the Transformer architecture and bipartite matching. While DETR streamlines the detection process, it requires 500 epochs to converge [93]. This spurs a wave of research that focuses on improving low training efficacy, including enhancing training signal [5; 15; 19; 94], the adoption of anchor boxes [49; 52], and the leverage of efficient attention mechanism [93; 11]. For instance, Co-DETR [94] introduces extra groups of object queries and DN-DETR [31] utilizes noisy queries to increase positive samples.
@@ 3 Hydra-SGG
§3.1 introduces our baseline model, which employs One-to-One Relation Assignment. Next, §3.2 describes Vanilla Hydra-SGG, which enhances the baseline by introducing a Hybrid Relation Assignment strategy. Finally, §3.3 details the complete Hydra-SGG model, which includes Hydra Branch (HydraBranch), an auxiliary decoder that promotes One-to-Many Relation Assignment. §3.4 performs a statistical analysis to validate our design of Hydra Branch, demonstrating its effectiveness in enhancing the One-to-Many assignment strategy.
@@ 3.1 One-to-One Relation Assignment Baseline
Baseline Architecture. Our baseline model is composed of a Backbone model, an Encoder, and a RelDecoder (Relation Decoder). The relation queries interact with image tokens extracted from the input image by Backbone and Encoder. The updated queries are processed by box regression, entity classification, and relation classification heads to generate the final predictions.
@@
  • One-to-One Relation Assignment Loss. In One-to-One Relation Assignment Loss $\mathcal{L}_{\text{o2o}}$ [7; 40], Hungarian matching (HM) is employed to find the optimal correspondence between the predicted relation triplets $\bar{\bm{R}}$ and the ground truth triplets $\bm{R}$. The loss function can be formulated as:
@@
    The classification losses (including entity and relation) and box regression losses are computed between matched predictions and ground truth labels, which is identical with the previous models [7; 40]. The Hungarian matching hyperparameters are the same as RelTR [7]. A detailed implementation of the loss function, training strategy, and model architecture is provided in Appendix. The evaluation of our baseline model is given in §4.3.
@@ 3.2 Vanilla Hybrid Relation Assignment
VG150 [75] contains an average of only 5.5 ground truth relation triplets per image in train. Our baseline adopts a One-to-One Relation Assignment and only assigns approximately 5 out of 300 relation queries to match the ground truth relation triplets for each image (i.e., $N\!=\!300$). This severe sparse supervision reduces the training efficacy, as only about 2% of queries in each training step match with ground truth labels for learning, while the remaining 98% are treated as negative samples. Consequently, the model requires a significantly higher number of training steps to learn effectively, leading to slow convergence.
@@
Figure 4: (a) The average number of positive samples of VG150 train and val for One-to-One and Hybrid Relation Assignment. (b) The percentage increase in positive samples achieved by Hybrid Relation Assignment compared to the One-to-One baseline. (c) ADS on VG150 val. (d)-(e) The visualizations show that for the same group of queries that previously predicted different relations, removing the self-attention layers causes them to make identical predictions. The Q ID column represents the ID of each relation query.
@@ 3.3 Complete Hydra-SGG
Each RelDecoder layer contains self- and cross-attention layers, and a point-wise feed-forward network (FFN). The self-attention layer enables inter-query interactions [69], while cross-attention and FFN do not explicitly support query interactions. Previous studies indicate that self-attention helps inhibit duplicate predictions [2; 7]. In this work, we further find that the self-attention layer in RelDecoder is crucial to reduce duplicated relation predictions.
@@
3.4 Statistical Analysis of Hydra Branch
To quantify how Hydra Branch enhances the One-to-Many assignment strategy, we analyze the Euclidean distances between query embedding pairs. Using 5,000 images from the VG150 val set, we compute pairwise distances for all 300 queries per image (totaling $\binom{300}{2}=44,850$ pairs per image) in two conditions: with and without self-attention layers. The removal of self-attention reduces the mean query distance from 7.55 to 7.23 (5% decrease). A paired t-test confirms statistical significance ($t=54.502$, $p<0.001$), with a Cohen’s $d$ effect size of 0.771 – approaching the threshold for a large effect ($d=0.8$). These results demonstrate that Hydra Branch promotes query similarity through self-attention removal, thereby facilitating One-to-Many assignment.